HETAV PATEL

Grade 11,

If the circumference of a circular sheet is 154 m, find its radius. Also find the area of the sheet. (Take π = 22/7)

 If the circumference of a circular sheet is 154 m, find its radius. Also find the area of the sheet. (Take π = 22/7)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,
Circumference of the circle = 154 m
Then,
We know that, Circumference of the circle = 2πr
154 = 2 × (22/7) × r
154= 44/7 × r
r = (154 × 7)/44
r = (14 × 7)/4
r = (7 × 7)/2
r = 49/2
r = 24.5 m
Now,
Area of the circle =πr^2
= 22/7 × (24.5)^2
= 22/7 × 600.25
= 22 × 85.75
= 1886.5 m^2
So, the radius of circle is 24.5 and area of circle is 1886.5.

Thanks